OpenStudy (rainbow_dash):

Find the LCM of the set of polynomials:

OpenStudy (rainbow_dash):

\[41a^3c, 8b^4, b^2c^2\]

OpenStudy (cwrw238):

do you know LCM of 41 and 8?

OpenStudy (cwrw238):

note 41 is a prime number

OpenStudy (rainbow_dash):

So it would be 41*8 then, right?? which is 328

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):

how about the letters?

OpenStudy (cwrw238):

b^2 and b^4 ?

OpenStudy (rainbow_dash):

well a would still be a^3 because there's no other a...the b's would be b^4 in the final result? and c is c^2 as well

OpenStudy (cwrw238):

absolutely right

OpenStudy (rainbow_dash):

awesome sauce-om. thank you :)

OpenStudy (cwrw238):

lol - yw

OpenStudy (rainbow_dash):

one quick question though. is the LCM the number you get when you take the lowest number that multiplies into each number you're given? or is it the lowest number the numbers you're given multiply to be

OpenStudy (cwrw238):

its the lowest number into which the numbers you are considering divide into if the numbers are large the best way to find the lcm is to write the numbers in prime factor form eg LCM of 156 and 180: 156 = 2*2 *3* 13 180 = 2*2*3*3*5 for each common factor take only one and use all the rest so LCM is 2*2*3*3*5*13

OpenStudy (rainbow_dash):

hmmm okay. thanks again

OpenStudy (cwrw238):

yw